A wire fixed at the upper end stretches by length ℓ by applying a force F. The work done in stretching is
(a) 2F ℓ
(b) f ℓ
(c) F/2 ℓ
(d) F ℓ/2

Grade:upto college level

3 Answers

Navjyot Kalra
9 years ago
(d). Work done by constant force in displacing the object by a distatance ℓ
= 1/2 * stress * strain * volume
= 1/2 / F/A * ℓ/L * A * L = Fℓ/2
Mihir
6 years ago
Work done is equal to 1/2 × stress × strain × volume U=1/2 × F/A × ∆l/l × AL = 1/2 × f ∆l ATQ = 1/2FL
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
Work done by constant force in displacing the object by a distatance ℓ
= 1/2 * stress * strain * volume
= 1/2 / F/A * ℓ/L * A * L = Fℓ/2
 
Thanks and regards,
Kushagra